# Recurrence Criteria for Reducible Homogeneous Open Quantum Walks on the Line

### v3

@@ 3.3 General Recurrence Criteria for HOQWs induced by a coin $(L,B,R)$ of dimension 2
Building upon the established preceding proposition, we will advance to the following theorem, delineating a comprehensive criterion for 2-dimensional coins, extending [15, Theorem 17] to encompass the Lazy case. This criterion will be structured into two distinct parts: the first adhering to the logic of the previous theorem, and the second articulated in terms of the eigenvalues of the operators of the coin, independent of the invariant densities of the auxiliary map.
@@
Now we suppose that $L,B,R$ share two linearly independent eigenvectors, then we can apply Theorem 3.9 to obtain $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ common orthonormal eigenvectors of $L,B,R.$ We assume that $|l_{1}|=|r_{1}|$ and $|l_{2}|=|r_{2}|,$ then for $\rho=\lvert u\rangle\langle u\rvert$, where $\lvert u\rangle$ is some common eigenvector of $L,B,R,$ we have $p_{j,j+1;\rho}(n)=p_{j,j-1;\rho}(n)=w$ for all $j\in\mathbb{Z}$ and all $n\geq 0$ for some constant $w$ and therefore, in this case, we have a classical homogeneous walk which is recurrent. If $(L,B,R)$ was not transient, then we would be in third situation of proposition 3.5, with
@@
4 A General Recurrence Criteria for Open Quantum Walks
Following the approach described in [10, Section 4], as well as the references therein, we can obtain for a coin $(L,R)$ a decomposition
@@ 4 A General Recurrence Criteria for Open Quantum Walks
Since $\mathfrak{h}$ is finite-dimensional, any minimal enclosure is included in $\mathfrak{R},$ and is the support of a unique extremal invariant state. Moreover, we have a unique decomposition of $\mathfrak{R}$ of the form
$\mathfrak{R}=\bigoplus_{\alpha\in A}\mathfrak{X}_{\alpha},$
@@
Further, every $\mathfrak{X}_{\alpha}$ either is a minimal enclosure or assumes a decomposition (not unique) as the sum of mutually orthogonal isomorphic minimal enclosures $\{\mathcal{Y}_{\alpha,\beta},\;\beta\in I_{\alpha}\}:$
@@
Given a minimal enclosure $\mathcal{Y}_{\beta},$ let $\tau_{\beta}$ be the unique invariant state with support on $\mathcal{Y}_{\beta}.$ Introducing the parameter
@@
we have by [10, Lemma 4.2] that $m_{\beta}$ is independent of the particular minimal enclosure $\mathcal{Y}_{\beta}$ in $\mathfrak{X}_{\alpha}.$ Therefore, we can combine Proposition 3.5, Theorem 3.8 and the discussion above to state the property
@@
  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)=\dfrac{1}{2}\Rightarrow$ the walk is $\tau_{\alpha}$-recurrent for all $\tau\in\mathcal{D}(\mathcal{Y}_{\alpha})$;
  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)\neq\dfrac{1}{2}\Rightarrow$ the walk is $\tau_{\alpha}$-transient for all $\tau\in\mathcal{D}(\mathcal{Y}_{\alpha}).$
Once the decomposition on Equation (5) is obtained, we can say that $\mathcal{Y}_{\alpha}$ is either recurrent or transient.
The following result is due to [10] and will be crucial to our last criterion. To state it, consider a coin $(L,R),$ and let $\mathcal{Y}$ be an enclosure for $\mathcal{L}.$ Denoting $p_{\mathcal{Y}}$ the orthogonal projection onto $\mathcal{Y},$ we can define the associated absorption operator (see [9]):
@@
Consider a coin $(L,R)$ with the decomposition of $\mathfrak{h}$ given on Equation (5), and let $a_{\alpha}(\rho)=\mathbb{E}_{\rho}\left[\mathrm{Tr}\left(A(\mathcal{Y}_{\alpha})\rho_{0}\right)\right].$ Then we have the convergence in law
@@
We remark that the convergence in law brought by corollary 4.2 can be upgraded to a convergence a.s., as demonstrated in [13, Theorem 3.5.2].
@@
5 Examples - 1D
@@
    It is worth noting that if we assume $m=0,$ then the walk is transient by Item (2) of Theorem 3.10. More explicitly, in this case, for the item (2.2), $|l_{1}|=0,\;|l_{2}|=n,\;|r_{1}|=n,\;|r_{2}|=0.$
@@
There is a unique invariant state $\rho_{\infty}$ given by
@@
Then
$\rho_{\infty}=\frac{1}{3}\begin{bmatrix}1&0\\ 0&2\end{bmatrix}$,
$\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)=3/7$ and $\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right)=5/21.$ The coin $(L,B,R)$ is transient. However, if we take
@@
then the invariant state is
@@
and $\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)=1/\sqrt{7}=\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right).$ Thus the coin $(L,B^{\prime},R)$ is recurrent.
@@
We remark that the choices of the coins above were taken in such a way that $L,B^{\prime}$ and $R$ had a common invariant eigenvector, namely $1,0]^{T},$ which was not an eigenvector of $B.$ This is due to the hard job of, given non-trivial $R$ and $L,$ to find $B^{\prime}$ in such a way that $(L,B^{\prime},R)$ is recurrent. Attempt were conduct using mathematic software with many $L$’s and $R$’s having simple entries, but even though the entries of $B$’s that made $(L,B,R)$ to be transient were very long to write in this paper.
@@
Let us consider the coin given in [10, Section 6.2]:

### v4

@@ 5.1 General Recurrence Criterion for HOQWs Induced by a Coin $(L,B,R)$ of Dimension 2
Building upon the established preceding theorem, we will advance to the following result, delineating a comprehensive criterion for 2-dimensional coins, extending [17, Theorem 17] to encompass the Lazy HOQW case. This criterion will be structured into two distinct parts: the first adhering to the logic of the previous theorem, and the second articulated in terms of the eigenvalues of the operators of the coin, independent of the invariant densities of the auxiliary map.
@@
Now we suppose that $L,B,R$ share two unit linearly independent eigenvectors $\lvert u_{1}\rangle,\lvert u_{2}\rangle.\;$ It is straightforward that $\rho_{1}=\lvert u_{1}\rangle\langle u_{1}\rvert$ and $\rho_{2}=\lvert u_{2}\rangle\langle u_{2}\rvert$ are distinct invariant states of the auxiliary map $\mathcal{L}$, thus $\mathcal{L}$ can not be ergodic. By Theorem 5.1, it has two invariant subspaces $S_{1}\neq\{0\}$ and $S_{2}\neq\{0\}$ such that $S_{1}\cap S_{2}=\{0\}.$ Since the coin has dimension 2, two distinct pure invariant states must correspond to orthogonal rank-1 projections. Indeed, if $\rho_{1}$ and $\rho_{2}$ are both invariant and $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ are linearly independent, the only way for $\mathcal{L}$ to have both as fixed points is if they are orthogonal. Otherwise, any operator that fixes both $\rho_{1}$ and $\rho_{2}$ would fix all operators, forcing $\mathcal{L}$ to be the trivial identity map. We assume that $|l_{1}|=|r_{1}|$ and $|l_{2}|=|r_{2}|,$ then for $\rho=\lvert u\rangle\langle u\rvert$, where $\lvert u\rangle$ is some common eigenvector of $L,B,R,$ we have $p_{j,j+1;\rho}(n)=p_{j,j-1;\rho}(n)=w$ for all $j\in\mathbb{Z}$ and all $n\geq 0$ for some constant $w$ and therefore, in this case, we have a classical homogeneous walk which is recurrent. If $(L,B,R)$ was not recurrent, then we would be in third situation of proposition 3.5, with
@@
5.2 General Recurrence Criterion for HOQWs Induced by Finite Dimensional Coins on the Line
Following the approach of [11, Section 4] and the references therein, we can obtain for a coin $(L,R)$ a decomposition
@@ 5.2 General Recurrence Criterion for HOQWs Induced by Finite Dimensional Coins on the Line
Since $\mathfrak{h}$ is finite-dimensional, any minimal enclosure is contained in $\mathfrak{R},$ and is the support of a unique extremal invariant state. Moreover, we have a unique decomposition of $\mathfrak{R}$ of the form
$\mathfrak{R}=\bigoplus_{\alpha\in A}\mathfrak{X}_{\alpha},$ (16)
@@
Further, every $\mathfrak{X}_{\alpha}$ is either a minimal enclosure or admits a decomposition (not unique) as the sum of mutually orthogonal isomorphic minimal enclosures $\{\mathcal{Y}_{\alpha,\beta},\;\beta\in I_{\alpha}\}:$
@@
Given a minimal enclosure $\mathcal{Y}_{\beta},$ let $\tau_{\beta}$ be the unique invariant state with support on $\mathcal{Y}_{\beta}.$ We introduce the parameter
@@
where $\tau_{\beta}$ is the unique invariant state supported on the minimal enclosure $\mathcal{Y}_{\beta}$.
Lemma 5.3.
[11, Lemma 4.2] The parameter $m_{\beta}$ is independent of the particular minimal enclosure $\mathcal{Y}_{\beta}$ in $\mathfrak{X}_{\alpha}$.
Consequently, $m_{\beta}$ is a characteristic parameter of the component $\mathfrak{X}_{\alpha}$ itself, and we may denote it by $m_{\alpha}$ when referring to the whole component. Therefore, we can combine Proposition 3.5, Theorem 3.8 and the discussion above to state the property
@@
  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)=\dfrac{1}{2}\Rightarrow$ the walk is $\tau$-recurrent for all $\tau\in\mathcal{D}(\mathcal{Y}_{\alpha})$;
  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)\neq\dfrac{1}{2}\Rightarrow$ the walk is $\tau$-transient for all $\tau\in\mathcal{D}(\mathcal{Y}_{\alpha}).$
Once the decomposition in Equation (17) is obtained, we can say that $\mathcal{Y}_{\alpha}$ is either recurrent or transient.
The following result is due to [11] and will be crucial to our next criterion. To state it, consider a coin $(L,R),$ and let $\mathcal{Y}$ be an enclosure for $\mathcal{L}.$ Denoting $p_{\mathcal{Y}}$ the orthogonal projection onto $\mathcal{Y},$ we can define the associated absorption operator (see [10]):
@@
[11, Corollary 4.5] Consider a coin $(L,R)$ with the decomposition of $\mathfrak{h}$ given in Equation (17), and let $a_{\alpha}(\rho)=\mathbb{E}_{\rho}\left[\mathrm{Tr}\left(A(\mathcal{Y}_{\alpha})\rho_{0}\right)\right].$ Then we have the convergence in law
@@
We remark that the convergence in law brought by Corollary 5.5 can be upgraded to a convergence a.s., as demonstrated in [15, Theorem 3.5.2].
@@
5.3 A Generalized Recurrence Criterion for HOQWs on the Grid
We have already presented some recurrence criteria for discrete-time and continuous-time OQWs on the set of vertices $\mathbb{Z}^{2}$, constrained by a minimal enclosure. These criteria are interesting tools for understanding the long-term behavior of OQWs. Building on these consequences, we aim to derive a stronger recurrence criterion for discrete-time HOQWs, allowing for a more comprehensive analysis of their statistics.
At this point we will allow the auxiliary map of a coin $(D)$ to have more than one invariant state, thus we consider the decomposition of $\mathfrak{h}$ given on Section 5.2. Therefore, given a minimal enclosure $\mathcal{Y}_{\alpha},$ let $\tau_{\alpha}$ be the unique invariant state with support on $\mathcal{Y}_{\alpha}.$ Consider the parameter
$\tilde{m}_{\alpha}=\begin{bmatrix}m_{1,\alpha}\\ m_{2,\alpha}\end{bmatrix},\quad m_{1,\alpha}=\mathrm{Tr}\left(D_{1}\tau_{\alpha}D_{1}^{*}\right)-\mathrm{Tr}\left(D_{3}\tau_{\alpha}D_{3}^{*}\right),\quad m_{2,\alpha}=\mathrm{Tr}\left(D_{2}\tau_{\alpha}D_{2}^{*}\right)-\mathrm{Tr}\left(D_{4}\tau_{\alpha}D_{4}^{*}\right).$
As stated in Lemma 5.3, the value $\tilde{m}_{\alpha}$ is independent of the particular minimal enclosure $\mathcal{Y}_{\alpha}$ in $\mathfrak{X}_{\alpha}.$ Therefore, we can combine Theorem 4.15 and results concerning CTOQWs above to state the next property, which is a continuous-time version on the grid of Proposition 5.4.
Proposition 5.8.
Consider a coin $(D).$ We have a decomposition
  • $\tilde{m}_{\alpha}=0\Rightarrow$ the walk is $\tau_{\alpha}$-recurrent for all $\tau\in\mathcal{D}(\mathcal{Y}_{\alpha})$;
  • $\tilde{m}_{\alpha}\neq 0\Rightarrow$ the walk is $\tau_{\alpha}$-transient for all $\tau\in\mathcal{D}(\mathcal{Y}_{\alpha}).$
Proof.
Fixed a subspace $\mathcal{Y}_{\alpha}$ which is the support of exactly one invariant state $\tau_{\alpha}\in\mathcal{D}(\mathfrak{h}),$ we assume that $\tilde{m}_{\alpha}=0.$ Then $(D)$ is $\rho$-recurrent for every $\rho$ by Theorem 4.11. If $\tilde{m}_{\alpha}\neq 0,$ then $(D)$ is transient by Theorem 4.12.
$\square$
With the decomposition in Equation (19) established, we can say that $\mathcal{Y}_{\alpha}$ is either recurrent or transient. It remains to understand what happens when a quantum state has support outside of $\oplus_{\alpha}\mathcal{Y}_{\alpha}.$
Theorem 5.9 (A Generalized Criterion for coins $(D)$).
[…]
For quick reference, Table 2 gathers the notation for the density sets and operator spaces employed in the paper.
Notation Description
$\mathfrak{h}$ Internal space
$\mathcal{D}(\mathfrak{h})$ Set of density operators on the internal space $\mathfrak{h}$
$\mathcal{D}_{i}(\mathfrak{h})$ Set of density operators on the internal space $\mathfrak{h}_{i}$
$\mathcal{S}(\mathcal{H}\otimes\mathcal{K})$ Set of density operators on $\mathcal{H}\otimes\mathcal{K}$ of diagonal form with respect to position
$\mathcal{D}(\mathcal{Y})$ Set of density operators supported in a subspace $\mathcal{Y}\subset\mathfrak{h}$
$\mathcal{D}_{1}$, $\mathcal{D}_{2}$ Sets of transient densities in Theorems 5.6 and 5.9
$\mathcal{I}_{1}(\mathcal{H})$ Trace-class operators on $\mathcal{H}$
$\mathcal{B}(\mathcal{W})$ Bounded linear operators on a Hilbert space $\mathcal{W}$
Table 2: Notation for operator sets and density spaces.
Having established rigorous recurrence criteria for HOQWs and CTOQWs on the line and on the grid, which classify the behavior based on the decomposition of the internal space $\mathfrak{h}$ and the drift parameters $\tilde{m}_{\alpha}$, we now illustrate these concepts with explicit numerical matrices. The following examples are also designed to demonstrate how the decomposition into enclosures are performed in practice.
6 Examples
6.1 1D
@@ 6 Examples
    Note that for $m=0$, the operators $L,B,R$ have the common and linearly independent eigenvectors $1,0]^{T}$ and $0,1]^{T}$. According to item (2.2) of Theorem 5.2, with $|l_{1}|=0$, $|l_{2}|=n$, $|r_{1}|=n$, $|r_{2}|=0$, the walk is transient.
@@
There is a unique invariant state $\rho_{\infty}$ for the auxiliary map given by
@@
Again, there is a unique invariant state $\rho_{\infty}$ for the auxiliary map $\mathcal{L}$ given by
$\rho_{\infty}=\frac{1}{3}\begin{bmatrix}1&0\\ 0&2\end{bmatrix}$.
In this case, $\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)=3/7$ and $\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right)=5/21,$ thus the coin $(L,B,R)$ is transient by item (1) of Theorem 5.2. However, if we take
@@
then the unique invariant state for the auxiliary map $\mathcal{L}^{\prime}$ is
@@
and $\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)=1/\sqrt{7}=\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right).$ Thus the coin $(L,B^{\prime},R)$ is recurrent, also by item (1) of Theorem 5.2
@@
We remark that this choice was possible because $L$ and $R$ share exactly one common eigenvector, namely $\lvert v\rangle=[1,0]^{T}$, and satisfy $L\lvert v\rangle=R\lvert v\rangle$. Under these conditions, $\mathrm{Tr}(L\rho L^{*})=\mathrm{Tr}(R\rho R^{*})$ holds for $\rho=\lvert v\rangle\langle v\rvert$. Hence, recurrence could be achieved by finding a matrix $B^{\prime}$ such that:
  1. 1.
    $\lvert v\rangle$ is also an eigenvector of $B^{\prime},$ thus $\rho$ becomes the unique invariant state for $\mathcal{L}^{\prime}$;
  2. 2.
    the normalization condition $L^{*}L+B^{\prime*}B^{\prime}+R^{*}R=I$ is satisfied.
The explicit matrix $B^{\prime}$ presented in the example was obtained by solving these constraints. It is worth noting that finding such $B^{\prime}$ for more general $L$ and $R$ is considerably more difficult. Several attempts were made using mathematical software (Maple) with various choices of $L$ and $R$ having simple entries; however, the resulting matrices $B^{\prime}$ that yielded recurrence often had very lengthy expressions, making them impractical to display in this paper.
Another natural continuation of this work would be to investigate under what conditions on $L$ and $R$ there exists a matrix $B^{\prime}$ such that $(L,B^{\prime},R)$ is recurrent, and to develop methods for explicitly constructing such $B^{\prime}$. Given fixed $L,R$, one would like to determine whether the set of admissible $B$ (satisfying $L^{*}L+B^{*}B+R^{*}R=I$) contains recurrent elements, and to describe them explicitly. As seen in this work, this problem is closely tied to the invariant state $\rho_{\infty}$ of $\mathcal{L}$ and the condition $\mathrm{Tr}(L\rho_{\infty}L^{*})=\mathrm{Tr}(R\rho_{\infty}R^{*})$.
@@
Let us consider the coin given in [11, Section 6.2]: